OpenStudy (korosh23):

Physics 12 Question! Satellites and Gravitation!

OpenStudy (korosh23):

For part a : I used g= GM/r^2 = 3.61 N/kg For part b: I don't know what is the force, please explain and show your work. Thank you

OpenStudy (paxpolaris):



OpenStudy (shamim):

F=mg+ma=?

OpenStudy (paxpolaris):

Net Force = Thrust Force + Weight and since we know the net force from: \[F_{net} = m \times a_{net}\] \[F_{net} = 87.5\ kg \times 1.20\ m/s^2 = \color{green}{105\ N}\] so, \[\Large 105 = x-316\]

OpenStudy (paxpolaris):

So the thrust force must be 421 N ... to counteract the gravity and accelerate upwards at 1.2 m/s^2

OpenStudy (priyar):

to normally lift an object we require mg upwards..but this needs to have an acc. of 1.2 so here the net acceleration is g+a u have already found out the 'g ' of mars as 3.61.. and mass and acceleration is already given..so put those values in our eq:F=m(g+a)..thats it!